OpenStudy (anonymous):

please help!! Find the slope of a line parallel 3x+5y=15. 5/3 3/5 - 5/3 - 3/5

OpenStudy (anonymous):

you can do this either by 1) solving for \(y\) or 2) memorize that the slope of \(ax=by=c\) is \(m=-\frac{a}{b}\)

OpenStudy (anonymous):

typo there, i meant \(ax+by=c\)

OpenStudy (anonymous):

huh?? im lost

OpenStudy (anonymous):

if you can turn \[3x+5y=15\] in to \[y=mx+b\] then the slope is \(m\) you do this by solving \(3x+5y=15\) for \(y\) do you know how to do that?

OpenStudy (anonymous):

nope still confuse

Directrix (directrix):

solve 3x+5y=15 for y First Step: Isolate the 5y term by subtracting 3x from both sides of the equation. 3x+5y=15 -3x -3x -------------- 5y = -3x + 15 Next step: divide both sides of the equation by 5. @knowel

OpenStudy (compassionate):

Parallel lines have the same slope. So solve and take this slope and, BAM!

OpenStudy (shamim):

3x+5y=15

OpenStudy (shamim):

5y=15-3x